OpenStudy (anonymous):

find the derivative of x 3root of x

OpenStudy (anonymous):

\[x \sqrt[3]{x}\]

hartnn (hartnn):

\(\huge x\sqrt[3]x=x^1.x^{1/3}=x^{4/3}\) now apply x^n formula of derivatives....u know that ?

OpenStudy (anonymous):

oh... okay 4/3x^1/3

hartnn (hartnn):

yes, thats correct! :)

OpenStudy (anonymous):

muahaha. thank you good sir!!

hartnn (hartnn):

welcome :)